What is the distance from point (-a,b) to point (2a,4b)? Thanks for your help.
Answer by TimothyLamb(4379)

Cartesian distance = d = sqrt( dx^2 + dy^2 )
---
dx = -a - 2a
dx = -3a
---
dy = b - 4b
dy = -3b
---
d = sqrt( (-3a * -3a) + (-3b * -3b) )
---
d = sqrt( 9aa + 9bb )
d = sqrt( 9(aa + bb) )
---
answer:
d = 3 * sqrt( aa + bb )
---
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php